A and B are running on a track having 8 laps to mile. If they run in opposite directions, they meet every 20 seconds. If they run in the same direction, it requires 3 minutes for A to overtake B. Find the rate of each in feet per second.


1
Expert's answer
2022-04-05T13:49:48-0400

1 mile is equal to 5280 feet, 1 lap is 5280/8=660 feet.

3 minutes are equal to 3"\cdot"60=180 seconds.

Let x - the rate of A in feet per second, y - the rate of B in feet per second.


"\begin{cases}
 (x+y) \cdot20=660, \\
 (x-y) \cdot180=660
\end{cases}"


"\begin{cases}
 x+y=33,\\
 x-y=3\cfrac{2}{3}
\end{cases}"


"\begin{cases}
 2x=36\cfrac{2}{3},\\
 y=33-x
\end{cases}"


"x =18\cfrac{1}{3},\ \ y =14\cfrac{2}{3} ."